Which of the following sets of quantum numbers is correct for an electron in 3d orbital?

A

N=3,L=2,M=3,S=+21

B

N=3,L=3,M=+3,S=21

C

N=3,L=2,M=2,S=+21

D

N=3,L=2,M=3,S=21

Grade:7

2 Answers

Geetha Reddy
3 years ago
Option c is correct. Set an electron in 3d orbital has quantum numbers N=3, l=2 , M=-2 ,S=+1/2. If n =3 then l=n-1=2 , if l=2 then m=-l to+l =-2to +2 , and s =+1/2 or -1/2
ankit singh
3 years ago

The following set of quantum numbers is correct for an electron in 3d orbital.
n=3,l=2,m=2,s=+21

For s,p,d and f orbitals, l=0,1,2,3 respectively.


When l=2, m can have values 2,1,0,+1,+2

The spin quantum number s can have value +21 or  21.